OpenStudy (anonymous):

Hello out there! I have a math question that I need an answer for. The question is..... When you have a number that has exactly two equal factors, each factor is called the S____ R____ of the number. It ask to fill in the blanks as well as give an example using the proper symbols. Thanks!

OpenStudy (anonymous):

is it like 4=2*2?

OpenStudy (anonymous):

So you have a number N with two equal factors = n. \[\implies N = n\times n = n^2 \implies n =\ ?\]

OpenStudy (anonymous):

What is n in terms of N.

OpenStudy (anonymous):

Square root.

OpenStudy (anonymous):

it says that N=n*n=n^2 Thus, \[n=\sqrt{N}\]